OpenStudy (buggiethebug):

A rock is thrown vertically with a speed of 62.8km/h. How long does it take to get to a height of 17m, if air resistance is ignored? I got 2.7x10^-4 hours, is that the answer?

OpenStudy (anonymous):

are the units correct for 62.8km/hr and 17m?

OpenStudy (buggiethebug):

yes

OpenStudy (buggiethebug):

i converted 17m to 0.017 km

OpenStudy (buggiethebug):

is my answer right?

OpenStudy (anonymous):

I got 3.34*10^-4hr. What did you use for g?

OpenStudy (buggiethebug):

I didnt really use any g

OpenStudy (buggiethebug):

V= 62.8km/h d= 0.017 km t=? V=d/t

OpenStudy (buggiethebug):

I was thinking that they didn't mention about the rock hitting the ground which means that it has nothing to do with acceleration due to gravity, they just simply asked when it reaches the height of 17m , they didn't mention where the rock is thrown from

OpenStudy (anonymous):

I would say you'd assume the rock is thrown from arbitrarily labeled x initial which=0. Either way. When you are dealing with the vertical component gravity is always considered.

OpenStudy (anonymous):



OpenStudy (anonymous):

vf^2=vi^2+2ad solve for vf. Now solve for t....d=t (vi+vf)/2